OpenStudy (anonymous):

Find the equation of the tangent plane to the surface z=16y2−0x2 at the point (0,−1,16). Find the equation of the tangent plane to the surface z=16y2−0x2 at the point (0,−1,16). @Mathematics

OpenStudy (anonymous):

take gradient of function z

OpenStudy (anonymous):

the gradient is vector normal to plane you are trying to find

OpenStudy (anonymous):

um....huh?

OpenStudy (anonymous):

ok, is that \[z=16y^2-0x^2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well you can ignore the x part since it is being multiplied by 0 so just z=16y^2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

rewrite like following 16y^2-z=0

OpenStudy (anonymous):

Take gradient \[{\partial f/dx,\partial f/dy,\partial f/dz}\] {0,32y,-1}

OpenStudy (anonymous):

k

OpenStudy (anonymous):

{0,32y,-1} is a vector normal to the plane you are trying to find plane equation v(x-a)+w(y-b)+x(z-c)

OpenStudy (anonymous):

yea...this is where i get lost

OpenStudy (anonymous):

i was with u..and i understand the formula here...but everytime i plug in the values..i get marked incorrect

OpenStudy (anonymous):

plug in this point at the point (0,−1,16) in here {0,32y,-1} plugged in numbers {0,-32,-1}

OpenStudy (anonymous):

v(x)+w(y+1)+t(z-16) 0(x)-32(y+1)-1(z-16)

OpenStudy (anonymous):

right so -32y-32-z+16

OpenStudy (anonymous):

simplifies to -32y-z=16

OpenStudy (anonymous):

is what i have.....so it's -z = 32y+32-16

OpenStudy (anonymous):

and z = -32y +48?

OpenStudy (anonymous):

i have to solve it for z =

OpenStudy (anonymous):

-32y-z=16 -z=32y+16 z=-32y-16

OpenStudy (anonymous):

yea see...says that's incorrect too

OpenStudy (anonymous):

lol...helps if i submit it to the correct question

OpenStudy (anonymous):

i was tryig to plug the answer into a different question....was correct

OpenStudy (anonymous):

you are in calc 3 right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but i missed an entire week due to being sick so i'm trying to teach myself this

OpenStudy (anonymous):

have you started multiple integral yet?

OpenStudy (anonymous):

no...we just took a test this past week...we're getting into the chain rule etc. next week

OpenStudy (anonymous):

oh, I have to review those because at my school , there is a calc 3 midterm and I am a tutor

OpenStudy (anonymous):

wow...ur school has tutor's for calc 3?

OpenStudy (anonymous):

no for all math.

OpenStudy (anonymous):

my school only goes up to calc 2 for tutors...and like 40 and hour for calc 3 tutors

OpenStudy (anonymous):

ours is free up to calc 2...then u gotta pay 40 an hour if you want tutoring for calc 3

OpenStudy (anonymous):

I should move to your school then ,

OpenStudy (anonymous):

yea really lol..where u go to school at?

OpenStudy (anonymous):

U of Illinois what about you?

OpenStudy (anonymous):

unc charlotte

OpenStudy (anonymous):

u of illinois is actually on my list to apply for grad schools...u like it up there?

OpenStudy (anonymous):

I like big school so it is great

OpenStudy (anonymous):

and middle of nowhere

OpenStudy (anonymous):

that's what i like

OpenStudy (anonymous):

wow, tutors at your school have it made

OpenStudy (anonymous):

yea...but we only get paid $9 an hour for other classes....it's until you get to the upper level classes where u actually make money

OpenStudy (anonymous):

which nobody wants to spend for so....

OpenStudy (anonymous):

I can help up to Linear algebra, and some logic courses

OpenStudy (anonymous):

i got 1 more of these i gotta create a new post for :)

OpenStudy (anonymous):

if u want some more medals

OpenStudy (anonymous):

sure why not